Question 1(Multiple Choice Worth 4 points)

(05.02)

Two quantities are related, as shown in the table:


xy

(2,3)

(4,4)

(6,5)

(8,6)


Which equation best represents the relationship? y = x + 2

 y = x + 1

 y = x + 2

 y = 2x + 1


1
Expert's answer
2021-11-02T10:06:16-0400

Slope "=m=\frac{Y_2-Y_1}{X_2-X_1}=\frac{4-3}{4-2}=\frac{5-4}{6-4}=\frac{6-5}{8-6}=\frac{1}{2}"


Equition of the line. "Y=mx-c"

"Y=\frac{x}{2}+C\\at\space X=2\\Y=3\\so,
3=\frac{2}{2}+C\\C=2"

"Y=\frac{x}{2}+2" will be the correct relationship.